OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x - 6y = 12 -2x + 3y = 6 -14 24 14 -24 so confused. help!

OpenStudy (anonymous):

please use an easy method.

OpenStudy (anonymous):

@svt2000

OpenStudy (anonymous):

im so sorry, i tried to to this,but couldnt.. let me give it another shot.

OpenStudy (anonymous):

ok thank you

OpenStudy (mertsj):

Multiply the bottom equation by 2

OpenStudy (mertsj):

What did you get?

OpenStudy (anonymous):

4x+6y=12 @ Mertsj

OpenStudy (mertsj):

Add that to the first equation.

OpenStudy (anonymous):

7x=24

OpenStudy (mertsj):

When you multiply the second equation by 2, you should get: -4x +6y = 12

OpenStudy (anonymous):

oh yeah I messed up on the - 2 so i'll get -x=24>>>>>x=-24

OpenStudy (mertsj):

yes. Now choose any equation that contains a y. Replace x with -24 and solve for y

OpenStudy (anonymous):

so......y=14

OpenStudy (mertsj):

Which equation did you use?

OpenStudy (anonymous):

3x - 6y = 12 3(-24)-6y=12 -72-6y=12 +72 -6y=84 84/-6 y=14

OpenStudy (mertsj):

When you divide + 84 by -6, what do you get?

OpenStudy (anonymous):

14

OpenStudy (mertsj):

Hmmm....I was always taught that positive divided by negative is negative.

OpenStudy (anonymous):

your right my bad

OpenStudy (anonymous):

-14

OpenStudy (mertsj):

That's what I got.

OpenStudy (anonymous):

thank you man your a life saver!

OpenStudy (mertsj):

your right my bad....that should be you're...short for you are and you are welcome (no extra charge for the English lesson)

OpenStudy (anonymous):

lol thank you anyway

OpenStudy (mertsj):

yw